OpenStudy (anonymous):

Please help with this calculus problem! The mass, in kilograms of salt, in a brine solution as a solution of different concentrations is being poured into the container is M = 20(1.5 + e^-0.02t), where t is the time in minutes. i) Determine an equation for the rate of change of the mass of salt in the brine solution. ii) After 5 mins, what is the rate of change of the mass of the salt in the solutions?

OpenStudy (callisto):

rate of change of M = dM/dt = d/dt [20(1.5 + e^-0.02t)] == d/dt [20(1.5) + 20e^-0.02t)] = 0+ 20(-0.02)e^-0.02t = -0.4e^-0.02t

OpenStudy (callisto):

for ii, put t=5 into dM/dt == -0.4e^-0.02t You'll get the rate of change of mass required :)

OpenStudy (anonymous):

WOW! You know you're the best right?

OpenStudy (callisto):

NO, i'm not!!!

OpenStudy (anonymous):

LOL, YES YOU ARE!!

OpenStudy (anonymous):

So, I substitute 5 into -0.4e^-0.02t?

OpenStudy (callisto):

Yes, it should be like that...

OpenStudy (anonymous):

yes he is

OpenStudy (callisto):

@JoBo Can you get the answer?

OpenStudy (anonymous):

It asks for the rate of change right, so I get -0.03619(what are the units?)

OpenStudy (anonymous):

kg/min?

OpenStudy (callisto):

Unit -> yes

OpenStudy (callisto):

answer -> should be yes

OpenStudy (anonymous):

LOL, THANKS :)

OpenStudy (callisto):

Welcome :)

OpenStudy (anonymous):

Im gonna post about 2 more, wanna help me on themÉ

OpenStudy (callisto):

I'll try ... I'm not sure if I know how to do it though

OpenStudy (anonymous):

lol ok Im pretty sure youère smart enough :P

OpenStudy (callisto):

Not really :(

OpenStudy (anonymous):

why are my question marks now ééés :(((

OpenStudy (callisto):

Hmmm.. I don't know too...

OpenStudy (anonymous):

K im gonna post a new one